40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3151     Entry time: Wed Jun 30 23:03:46 2010
Author: rana 
Type: Configuration 
Category: IOO 
Subject: Power into MC restored to max 

Kiwamu, Nancy, and I restored the power into the MC today:

  1. Changed the 2" dia. mirror ahead of the MC REFL RFPD back to the old R=10% mirror.
  2. Since the MC axis has changed, we had to redo the alignment of the optics in that area. Nearly all optics had to move by 1-2 cm.
  3. 2 of the main mounts there had the wrong handedness (e.g. the U100-A-LH instead of RH). We rotated them to some level of reasonableness.
  4. Tuned the penultimate waveplate on the PSL (ahead of the PBS) to maximize the transmission to the MC and to minimize the power in the PBS reject beam.
  5. MC_REFL DC  =1.8 V.
  6. Beams aligned on WFS.
  7. MC mirrors alignment tweaked to maximize transmission. IN the morning we will check the whole A2L centering again. If its OK, fine. Otherwise, we'll restore the bias values and align the PSL beam to the MC via the persicope.
  8. waveplates and PBS in the PSL were NOT removed.
  9. MC TRANS camera and QPD have to be recentered after we are happy with the MC axis.
  10. MC REFL camera has to be restored.
  11. WFS measurements will commence after the SURF reports are submitted.

We found many dis-assembled Allen Key sets. Do not do this! Return tools to their proper places or else you are just wasting everyone's time!

 

ELOG V3.1.3-